How many gallons does a 1'x 3' x 16" fish tank hold?
Answer by mananth(16949)

1'x 3' x 16"
convert the dimensions to feet
1' X 3/12' x 5/4'
=15/48
=5/16 cubic feet
..
1 cubic feet = 7.48 gallons
5/16 cubic feet = 7.48*5 /16
= 2.33 gallons